OpenStudy (anonymous):

write an equation for the line in slope-intercept form passing through (2, 3) and (8, 2)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ 3}})\quad &({\color{red}{ 8}}\quad ,&{\color{blue}{ 2}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (triciaal):

@jdoe0001 that looks nice in color @hahaalgebra do you have any questions after following the above?

OpenStudy (anonymous):

i cannot get the point slope form

OpenStudy (anonymous):

I came out with -1/6

OpenStudy (triciaal):

slope-intercept form y = m x + b where m is the slope and b is the y-intercept

OpenStudy (triciaal):

yes the slope is correct -1/6

OpenStudy (anonymous):

so is it y-3=-1/6(x-2)

OpenStudy (triciaal):

use one of the given points, plug the x, y and m then find b

OpenStudy (triciaal):

yes you have it in point-slope form and the question ask for slope-intercept form so now just rearrange to y =

OpenStudy (anonymous):

so I add 3 to both sides

OpenStudy (triciaal):

yes you also need to expand the parenthesis

OpenStudy (triciaal):

y = m x + b

OpenStudy (anonymous):

i am lost y = -1/6 + 10/3

OpenStudy (triciaal):

you lost your x

OpenStudy (anonymous):

oops y = -1/6x + 10/3

OpenStudy (triciaal):

y-3=-1/6(x-2) y - 3 = -1/6(x) -1/6(-2) y = -1/6 x + 10/3

OpenStudy (triciaal):

hint first check what format the answer should be in (standard, point-slope, or slope-intercept) for this one requested slope-intercept so just plug with the given point to find b of course you see you will get the same answer to put in point-slope and then rearrange

OpenStudy (triciaal):

and you know the slope is the change in y / corresponding change in x

OpenStudy (anonymous):

thank you for your help